OpenStudy (anonymous):

please explain c+(4-3c)+2=

OpenStudy (anonymous):

c+4-3c-2=0 -2c+2=0 -2c=-2 c=1 c+(4-3c)-2=0 Add 2 both sides: c+(4-3c)-2+2=2 Simplify -2+2 into 0: c+(4-3c)=2 Switch the order of the 4 and -3c (Commutative Property for Addition): c+(-3c+4)=2 Move the parentheses: [c+(-3c)]+4=2 Add c and -3c: -2c+4=2 Subtract 4 from both sides: -2c+4-4=2-4 Simplify 4-4 and 2-4: -2c=-2 Divide both sides by -2: c=1 Therefore, c=1.

OpenStudy (anonymous):

You drop down the c to the 2 understand?

OpenStudy (anonymous):

Wait Ha! I messed up don't pay attention to that

OpenStudy (anonymous):

Wait ... No I did not. I gived you the answer correct I just made it long

OpenStudy (anonymous):

Simplifying C + (4 + -3c) + -2 = 0 Remove parenthesis around (4 + -3c) C + 4 + -3c + -2 = 0 Reorder the terms: 4 + -2 + C + -3c = 0 Combine like terms: 4 + -2 = 2 2 + C + -3c = 0 Solving 2 + C + -3c = 0 Solving for variable 'C'. Move all terms containing C to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + C + -2 + -3c = 0 + -2 Reorder the terms: 2 + -2 + C + -3c = 0 + -2 Combine like terms: 2 + -2 = 0 0 + C + -3c = 0 + -2 C + -3c = 0 + -2 Combine like terms: 0 + -2 = -2 C + -3c = -2 Add '3c' to each side of the equation. C + -3c + 3c = -2 + 3c Combine like terms: -3c + 3c = 0 C + 0 = -2 + 3c C = -2 + 3c Simplifying C = -2 + 3c

OpenStudy (anonymous):

Is that clearly ?

OpenStudy (anonymous):

yes ma'am! thank you so much!

OpenStudy (anonymous):

Your very welcome (: